Stephen plans to purchase a car 7 years from now. The car will cost $50,052 at that time. Assume that Stephen can earn 7.53 percent (compounded monthly) on his money. How much should he set aside today for the purchase?

Round the answer to two decimal places.
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

r = 7.53%/12 = 0.6275% (monthly)

n = 7 * 12 = 84 months

$50,052 = PV * (1 + 0.006275)84

50,052 = PV * 1.691225

PV = $29,595.12